OpenStudy (vanz):

how do you solve this x2 + 3x - 9 = 0 ?

hero (hero):

You have to complete the square

OpenStudy (vanz):

i did, but i want to check my answer :)

OpenStudy (pokemon23):

Hero good news the test was easier than I thought If I pass I go to a regular math class :)

hero (hero):

x^2 + 3x = 9 x^2 + 3x + (3/2)^2= 36/4 + 9/4 (x+3/2)^2 = 45/4 x + 3/2 = +/- sqrt{45}/2 x = +/-3sqrt{5}/2 - 3/2 x = (+/- 3sqrt{5} - 3)/2